[figure: 9.7 cm]

THERMAL EXPANSION ANOMALIES AT A QUADRUPOLAR

ORDERING IN DyVO_{4}

Z.A. Kazei^{1}, N.P. Kolmakova^{2}, I.B. Krynetskii^{1}, A.A. Sidorenko^{2} and L.V. Takunov^{2} 1 Moscow State University, 119899 Moscow, Russia 2 Technical University, 241035 Bryansk, Russia

The thermal expansion anomalies in DyVO_{4} at a spontaneous ordering of the quadrupolar moments of Dy^{3+} ions are studied both experimentally, with the help of glued strain gauges and a capacitive deformation sensor, and theoretically. Characteristic anomalies in the unit cell parameters Δa/a(T), Δb/b(T) and Δc/c(T) as well as in the Δa′/a′ curve along the non active Jahn Teller direction are revealed for T < T_{c}=14 K. Within the general crystal-field approach the magnetoelastic contributions to Δa/a, Δb/b, Δc/c and Δa′/a′ both from the dominated Jahn-Teller modes ϵ^{γ} and totally symmetric modes ϵ^{α1} and ϵ^{α2} are calculated. The contributions of the totally symmetric modes, proportional to the quadrupole moment < O_{2}^{0} >, result in observable magnetoelastic effects, in particular in the tetragonal and volume deformations, and its calculation has required taking into account all the details of the crystal field. The values of magnetoelastic B^{γ} and total quadrupolar G^{γ} coefficients are determined, From the data on spontaneous deformations and spectroscopy the contribution of the pair quadrupolar interactions K^{γ} to the G^{γ} being positive. It is shown that the thermal expansion of DyVO_{4}, both in tetragonal and orthorhombic phases, is well described by a common model using a single set of interaction parameters. This work is supported by Russian Ministry of Education (grant N 97-0-7.3-157).

[figure: 13.4 cm]

Subject category :

5. Phase Transitions and Critical Phenomena

Presentation mode : poster

Corresponding author : I.B. Krynetskii

Address for correspondence : Department of Physics, Moscow State University, Vorob’evy Gory 119899, Moscow, Russia

Email address : krynets@plms.phys.msu.su